Tape is wound around a disk shaped object which is at rest On a frictionless surface. Assume the disk is homogeneous, has mass M and radius R. The free end of tape is pulled horizontally with a force F. Find an expression for horizontal distance d covered by the disk as a function of length l of the tape unwound. Ignore the mass the tape as compared with the mass of the disk.
1
Expert's answer
2020-07-13T11:41:21-0400

Moment of inertia of the solid disk:


"I=\frac{MR^2}2."


When we pull the tape, it creates a torque:


"\tau=FR."

This torque creates angular acceleration:


"\alpha=\frac{\tau}{I}=\frac{2F}{MR}"

This angular acceleration provides tangential acceleration:


"a_\tau=\alpha R=\frac{2F}{M}."


As the tape unwinds, the disk rotates for some angle "\theta". If the length of the tape unwound is L, the angle corresponds to


"\theta=\frac{L}{\pi}."

For the disk to unwind for angle "\theta" (from bottom to top are the surface, disk, and tape), it will take a time of


"t=\sqrt{\frac{2\theta}{\alpha}}=\sqrt{\frac{2L}{\alpha\pi}}=\sqrt{\frac{LMR}{F\pi}}."


Meanwhile under the influence of linear acceleration


"a_l=\frac{F}{M}"

the disk will travel distance d for the same time:


"t=\sqrt{\frac{2d}{a_l}}=\sqrt{\frac{2dM}{F}}."

Equate time expressed in terms of d and L:


"\sqrt{\frac{2dM}{F}}=\sqrt{\frac{LMR}{F\pi}},\\\space\\
d(L)=\frac{R}{2\pi}L."